A 3.35 g sample of an unknown gas at 37 ∘C and 1.05 atm is stored in a 2.75 L flask.

What is the density of the gas?

density:

g/L

What is the molar mass of the gas?

molar mass:
Answer #1

Molk mass of gar = 3.55(2), T = 37% = (37+2731k=310k P = 1.05 atm, Volume (V) = 2.75, L, R = 01082 liath Ideal gas law PV = n